OpenStudy (anonymous):

solve this equation 2x/3 - 1/2 = 2x+5/6

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

how you get that

OpenStudy (anonymous):

I'll show you, one sec...

OpenStudy (anonymous):

Are you going to show me :)

OpenStudy (anonymous):

2x/3 - 1/2 = 2x + 5/6 Add (1/2 - 2x) to both sides (-4x/3) = 4/3 Divide both sides by (-4/3) x = -1

OpenStudy (anonymous):

@luisrock2008 helloooo? does this help you? :P

OpenStudy (anonymous):

Yes.